rebeccaxhawaii (rebeccaxhawaii):

Given the functions f(n ) = 11 and g(n ) = -2(n - 1), combine them to create an arithmetic sequence, an, and solve for the 31st term.

ILovePuppiesLol (ilovepuppieslol):

noob, this is so easy

rebeccaxhawaii (rebeccaxhawaii):

says the one in alg 1 -.-

ILovePuppiesLol (ilovepuppieslol):

@Kainui help us please :)

OpenStudy (inkyvoyd):

noob, this is so easy

ILovePuppiesLol (ilovepuppieslol):

well said @inkyvoyd

ILovePuppiesLol (ilovepuppieslol):

@jim_thompson5910 @Mehek14 please help us :(

jimthompson5910 (jim_thompson5910):

one way is to do f(n) + g(n) = [11] + [-2(n-1)] f(n) + g(n) = 11 - 2(n-1) notice how 11 is the first term and -2 is the common difference

OpenStudy (kainui):

I have honestly no idea what you're expected to do

jimthompson5910 (jim_thompson5910):

you could also do f(n) - g(n), but that leads to a different common difference

rebeccaxhawaii (rebeccaxhawaii):

then plug 31 for n?

ILovePuppiesLol (ilovepuppieslol):

wow my tags worked

rebeccaxhawaii (rebeccaxhawaii):

yes ty b

jimthompson5910 (jim_thompson5910):

`then plug 31 for n?` yes correct @rebeccaxhawaii

rebeccaxhawaii (rebeccaxhawaii):

okay so 11-2(31-1) 11-2(30) 11-60 f(n)+g(n)=-49??

jimthompson5910 (jim_thompson5910):

looks good

rebeccaxhawaii (rebeccaxhawaii):

yay thank you c:

jimthompson5910 (jim_thompson5910):

no problem